# 3D-LFM: Lifting Foundation Model

### v2

@@ 3 Approach
We scale the zero-centered data to the range $[-1,1]$ while preserving the aspect ratio to maintain the geometric integrity of the keypoints. For more details on handling missing data in the presence of perspective effects, we refer the reader to Deep NRSFM++[25].
@@
where $D$ is the dimensionality of the Fourier feature space, $\boldsymbol{\omega}\in\mathbb{R}^{2\times\frac{D}{2}}$ and $\mathbf{b}\in\mathbb{R}^{\frac{D}{2}}$ are parameters sampled from a normal distribution, scaled appropriately. These parameters are sampled once and kept fixed, as per the RFF methodology. The output of this transformation $\mathbf{TPE}(\mathbf{W}_{c})$ is then fed into the graph-based transformer network as $\mathbf{X}^{\ell}$ where $\ell$ indicates the layer number (00 in the above case). This set of features is now ready for processing inside the graph-based transformer layers without the need for correspondence among the input keypoints. The TPE retains the property of permutation equivariance while implicitly encoding the relative positions of the keypoints.
3.1 Graph-based Transformer Architecture
@@
Hybrid Feature Aggregation: For each layer $\ell$, ranging from 0 to $L$, the feature matrix $\mathbf{X}^{(\ell)}\in\mathbb{R}^{N\times D}$ is augmented through simultaneous local and global processing. The local processing component, GA($\mathbf{X}^{(\ell)},\mathbf{A}$), leverages an adjacency matrix $\mathbf{A}$, which encodes the connectivity based on the object category, to perform graph-based attention on batches of nodes representing the input 2D data:
@@ 3.2 Procrustean Alignment
$\underset{\mathbf{R}}{\text{minimize}}\quad||\mathbf{M}\odot(\mathbf{S}_{r}-\mathbf{S}_{c}\mathbf{R})||_{F}^{2}$
@@ 3.3 Loss Function
4 Results and Comparative Analysis
Our evaluation shows the 3D Lifting Foundation Model (3D-LFM)’s capability in single-frame 2D-3D lifting across diverse object categories without object-specific data in Sec. 4.1. Following that, Sec. 4.2 highlights 3D-LFM’s performance over specialized methods, especially achieving state-of-the-art performance in whole-body benchmarks[32] (Fig. 3). Additionally, Sec. 4.3 shows 3D-LFM’s capability in 2D-3D lifting across 30 categories using a single unified model, enhancing category-specific performance and achieving out-of-distribution (OOD) generalization for unseen object configurations during training. In conclusion, the ablation studies in Section 4.4 validate our proposed procrustean approach, token positional encoding, and the local-global hybrid attention mechanism in the transformer model, confirming their role in 3D-LFM’s effectiveness in both single- and multiple-object scenarios.
4.1 Multi-Object 3D Reconstruction
@@
4.2 Benchmark: Object-Specific Models
@@
4.3 Towards foundation model
In this section, we highlight 3D-LFM’s role as a foundational model for varied 2D-3D lifting, capable in managing multiple object types and data imbalances. In this subsection, we explore 3D-LFM’s scalability for collective dataset training (Sec.4.3.1), its generalization to new categories and rig transfer capabilities (Sec.4.3.2). These studies validate the 3D-LFM’s role as a foundation model, capable at leveraging diverse data without requiring specific configurations, thus simplifying the 3D lifting process for varied joint setups.
@@ 4.3 Towards foundation model
4.3.1 Combined Dataset Training
@@ 4.3.2 OOD generalization and rig-transfer:
Additionally, we show 3D-LFM’s capability in transferring rig configurations between datasets, embodying the concept of generic geometry learning. By training on a $17$-joint Human3.6M dataset [7] and testing on a $15$-joint Panoptic Studio setup [9], our model gives accurate 3D reconstructions despite variations in joint arrangements. This capability is particularly interesting for its efficiency in utilizing data from multiple rigs of the same object, and underscores the model’s adaptability, a cornerstone in processing diverse human datasets. It aligns with the broader community’s interest in versatile geometry learning, which makes these findings especially compelling. For a more thorough validation, we direct the reader to the ablation section, where qualitative visuals (Fig. 6) and quantitative analysis (Sec. 4.4.3) further highlight 3D-LFM’s OOD generalization and rig transfer efficacy.
@@ 4.4 Ablation
4.4.1 Procrustean Transformation
3D-LFM’s fusion of the procrustean approach, a first in transformer-based lifting frameworks, concentrates on deformable object components, as outlined in Sec.3.2. By focusing on shape within a standard canonical reference frame and avoiding rigid body transformations, we see faster learning and a decreased MPJPE, as evident by the gap between blue and orange lines in Fig. 5 (a) suggests. This fusion is crucial for learning 3D deformations, while utilizing transformers’ equivariance. These findings suggest that even for transformers, avoiding rigid transformations’ learning aids convergence, most notably with imbalanced datasets.
@@ 4.4.1 Procrustean Transformation
4.4.2 Local-Global vs. Hybrid Attention
@@
4.4.3 Tokenized Positional Encoding:
@@
Data imbalance study: When tested on the underrepresented hippo category from the Animal3D dataset [27], TPE based model showed a $3.27$% improvement in MPJPE over the baseline MLP with a $128$-dimensional model performance as evident in first row of Tab. 3. This improvement grew with the model size. These results highlight TPE’s scalability and its faster convergence, especially relevant in imbalanced, OOD scenarios as detailed in Fig. 5 (b). The observed performance boosts suggest that TPE’s analytical nature might be more suited to adapting to novel data distributions. Increasing model size amplifies TPE’s benefits, hinting that its fixed analytical approach more adeptly handles OOD intricacies compared to learnable methods like MLPs, which may falter in such situations.
@@
Rig transfer study: Our rig transfer analysis, summarized in Table 3, showcases TPE’s adaptability and effectiveness over the MLP baseline across different joint configurations and rig scenarios, with improvements up to $52.3$%. These findings, particularly the significant performance boost in complex rig transfers, underscore TPE’s robustness in OOD contexts. Figure 6 visually highlights the qualitative differences between TPE and MLP approaches in a rig transfer scenario, where the model trained on a $17$-joint [7] configuration is tested on a $15$ joint [9] setup.
5 Discussion and Conclusion
@@
Acknowledgement: We extend our gratitude to Ian R. Fasel, Tim Clifford, Javier Movellan, and Matthias Hernandez of Apple for their insightful discussions.
@@ II Limitations
Figure 7: Challenges in Perspective and Depth Perception: (a) Incorrect leg orientation due to depth ambiguity in monkey capture. (b) Misinterpreted head position in a second monkey example. (c) A tiger’s keypoints distorted by perspective, leading to primate-like 3D predictions."

### v3

@@ III Approach
We scale the zero-centered data to the range [1,1][-1,1] while preserving the aspect ratio to maintain the geometric integrity of the keypoints. For more details on handling missing data in the presence of perspective effects, we refer the reader to Deep NRSFM++[24].
@@
where $D$ is the dimensionality of the Fourier feature space, $\boldsymbol{\omega}\in\mathbb{R}^{2\times\frac{D}{2}}$ and $\mathbf{b}\in\mathbb{R}^{\frac{D}{2}}$ are parameters sampled from a normal distribution, scaled appropriately. These parameters are sampled once and kept fixed, as per the RFF methodology. The output of this transformation $\mathbf{TPE}(\mathbf{W}_{c})$ is then fed into the graph-based transformer network as $\mathbf{X}^{\ell}$ where $\ell$ indicates the layer number (0 in the above case). This set of features is now ready for processing inside the graph-based transformer layers without the need for correspondence among the input keypoints. The TPE retains the property of permutation equivariance while implicitly encoding the relative positions of the keypoints.
III.1 Graph-based Transformer Architecture
@@
[…]
@@ III.2 Procrustean Alignment
[…]
@@ III.3 Loss Function
IV Results and Comparative Analysis
Our evaluation shows the 3D Lifting Foundation Model (3D-LFM)’s capability in single-frame 2D-3D lifting across diverse object categories without object-specific data in Sec. IV.1. Following that, Sec. IV.2 highlights 3D-LFM’s performance over specialized methods, especially achieving state-of-the-art performance in whole-body benchmarks[32] (Fig. 3). Additionally, Sec. IV.3 shows 3D-LFM’s capability in 2D-3D lifting across 30 categories using a single unified model, enhancing category-specific performance and achieving out-of-distribution (OOD) generalization for unseen object configurations during training. In conclusion, the ablation studies in Section IV.4 validate our proposed procrustean approach, token positional encoding, and the local-global hybrid attention mechanism in the transformer model, confirming their role in 3D-LFM’s effectiveness in both single- and multiple-object scenarios.
IV.1 Multi-Object 3D Reconstruction
@@
IV.2 Benchmark: Object-Specific Models
@@
IV.3 Towards foundation model
In this section, we highlight 3D-LFM’s role as a foundational model for varied 2D-3D lifting, capable in managing multiple object types and data imbalances. In this subsection, we explore 3D-LFM’s scalability for collective dataset training (Sec.IV.3.1), its generalization to new categories and rig transfer capabilities (Sec.IV.3.2). These studies validate the 3D-LFM’s role as a foundation model, capable at leveraging diverse data without requiring specific configurations, thus simplifying the 3D lifting process for varied joint setups.
@@ IV.3 Towards foundation model
IV.3.1 Combined Dataset Training
@@ IV.3.2 OOD generalization and rig-transfer:
Additionally, we show 3D-LFM’s capability in transferring rig configurations between datasets, embodying the concept of generic geometry learning. By training on a $17$-joint Human3.6M dataset [7] and testing on a $15$-joint Panoptic Studio setup [9], our model gives accurate 3D reconstructions despite variations in joint arrangements. This capability is particularly interesting for its efficiency in utilizing data from multiple rigs of the same object, and underscores the model’s adaptability, a cornerstone in processing diverse human datasets. It aligns with the broader community’s interest in versatile geometry learning, which makes these findings especially compelling. For a more thorough validation, we direct the reader to the ablation section, where qualitative visuals (Fig. 6) and quantitative analysis (Sec. IV.4.3) further highlight 3D-LFM’s OOD generalization and rig transfer efficacy.
@@ IV.4 Ablation
IV.4.1 Procrustean Transformation
3D-LFM’s fusion of the procrustean approach, a first in transformer-based lifting frameworks, concentrates on deformable object components, as outlined in Sec.III.2. By focusing on shape within a standard canonical reference frame and avoiding rigid body transformations, we see faster learning and a decreased MPJPE, as evident by the gap between blue and orange lines in Fig. 5 (a) suggests. This fusion is crucial for learning 3D deformations, while utilizing transformers’ equivariance. These findings suggest that even for transformers, avoiding rigid transformations’ learning aids convergence, most notably with imbalanced datasets.
@@ IV.4.1 Procrustean Transformation
IV.4.2 Local-Global vs. Hybrid Attention
@@
IV.4.3 Tokenized Positional Encoding:
@@
Data imbalance study: When tested on the underrepresented hippo category from the Animal3D dataset [27], TPE based model showed a 3.273.27% improvement in MPJPE over the baseline MLP with a $128$-dimensional model performance as evident in first row of Tab. 3. This improvement grew with the model size. These results highlight TPE’s scalability and its faster convergence, especially relevant in imbalanced, OOD scenarios as detailed in Fig. 5 (b). The observed performance boosts suggest that TPE’s analytical nature might be more suited to adapting to novel data distributions. Increasing model size amplifies TPE’s benefits, hinting that its fixed analytical approach more adeptly handles OOD intricacies compared to learnable methods like MLPs, which may falter in such situations.
@@
Rig transfer study: Our rig transfer analysis, summarized in Table 3, showcases TPE’s adaptability and effectiveness over the MLP baseline across different joint configurations and rig scenarios, with improvements up to 52.352.3%. These findings, particularly the significant performance boost in complex rig transfers, underscore TPE’s robustness in OOD contexts. Figure 6 visually highlights the qualitative differences between TPE and MLP approaches in a rig transfer scenario, where the model trained on a $17$-joint [7] configuration is tested on a $15$ joint [9] setup.
V Discussion and Conclusion
@@
Acknowledgements
This research was supported in part by the Australian Research Council under Discovery Project DP220103803.
@@ II Limitations
Figure 7: Challenges in Perspective and Depth Perception: (a) Incorrect leg orientation due to depth ambiguity in monkey capture. (b) Misinterpreted head position in a second monkey example. (c) A tiger’s keypoints distorted by perspective, leading to primate-like 3D predictions.”